OpenStudy (anonymous):

Latonya bought flowers for her mother. She spent $45. Each tulip was $1 and each rose was $2. She bought 3 more tulips than roses. Let t be the number of tulips. Which algebraic sentence would help you figure out how many tulips Latonya bought?

OpenStudy (anonymous):

A. t + 2(t + 3) = 45 B. t + 2(t – 3) = 45 C. 2t (t + 3) = 45 D. -3t + 2t = 45

OpenStudy (anonymous):

B